OpenStudy (anonymous):

Graph the first six terms of a sequence where a1=4 and r=2. Please help I am stumped. will award medal /points to best answer

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[a_1,a_2,a_3,a_4,a_5\\ a_1,a_1\times ra_1\times r^2,a_1\times r^3,a_1\times r^4\]

OpenStudy (anonymous):

\[4,4\times 2,4\times 2^2,4\times 2^3,4\times 2^4\]

OpenStudy (anonymous):

@satellite73 I dont understand can you please explain

OpenStudy (anonymous):

@rock_mit182

OpenStudy (rock_mit182):

given a sequence: a1, a2, a3 .... an; for n= 1,2, 3 . .. here each term after the first is found by multiplying the previous one by a fixed, non-zero number called the common ratio(r)

OpenStudy (rock_mit182):

The general form of a geometric sequence is a,\ ar,\ ar^2,\ ar^3,\ ar^4,\

OpenStudy (rock_mit182):

where r ≠ 0

OpenStudy (anonymous):

okay that makes sense so i would multiply my a1 term by my r term?

OpenStudy (rock_mit182):

yes but remember the next term is:\[a _{n+1} = a _{n}*r\]

OpenStudy (rock_mit182):